OpenStudy (anonymous):

In the standard (x,y) coordinate plane, what is the slope of the line that passes through the points (-4,3) and (6,-5)?

OpenStudy (anonymous):

-4/5

OpenStudy (anonymous):

(-5-3)/(6-4)= -8/2= -4 3=-4(-4)+b 3=16+b 3-16=b -13=b So, y=-4x-13

OpenStudy (anonymous):

slope = (y2-y1)/(x2-x1)

OpenStudy (anonymous):

I am not sure how you got -4/5